# Hierarchical Graph Neural Network for Compressed Speech Steganalysis

### v1

@@ 3.2 QIM-based Steganography
[…]
@@
The challenge in QIM steganography arises from increased distortion resulting from the use of a reduced codeword set in the quantization procedure. The fundamental issue lies in how to partition the original codebook $C_{i}$ into multiple sub-codebooks. Researchers have proposed various methods including matrix encoding QIM (ME-QIM) [tian2014improving], complementary neighbour vertices QIM (CNV-QIM) [xiao2008approach], and others.
@@ 3.3 GNN Overview
[…]
A widely adopted family of GNN models includes graph convolutional networks (GCNs) [kipf2016semi], gated graph neural networks (GGNNs) [li2015gated], graph sample and aggregation (GraphSAGE) [hamilton2017inductive], and more. These architectures often involve variations of the basic propagation rule to capture more complex relationships within the graph.
@@ 4 Proposed Framework
Building upon the foundational concepts of speech coding, QIM steganography, and GNN discussed in Section 3, we now present the detailed architecture and components of our proposed steganalysis system. Specifically, this section describes an efficient method designed to uncover QIM steganography within compressed, LBR VoIP streams. Our approach is specifically tailored to work with speech compressed using the G.729A codec, which is widely used in VoIP communications. Figure 2 illustrates the structure of our proposed VoIP steganalysis, while Algorithm 1 outlines the steps involved. The current section provides a detailed explanation of the proposed approach, which includes codewords correlation analysis, graph construction, and network architecture.
Figure 2: The overall structure of the proposed approach.
@@
13:Construct adjacency matrix $A$:
@@ 4.1 Codewords correlation analysis
Due to the repeating patterns in human speech sounds (especially vowels and voiced consonants), the vocal signal exhibits stability over limited durations, approximately the frame length. Consequently, the codewords exhibit correlations within the same frame. Furthermore, speech signals have local periodicity, meaning that codewords have similar values in different frames. As a result, there are four types of inter-codeword correlations, namely: intra-frame correlation, successive frame correlation, cross-frame correlation, and cross-word correlation [lin2018rnn]. Local features are characterized by the first two correlations, while global ones are reflected by the latter two.
The QIM steganography process has a direct impact on the correlations between these codewords, thereby altering their statistical distribution. Earlier VoIP steganalysis approaches aimed to extract correlation features to uncover steganograms, as in [li2017steganalysis, yang2018steganalysis].
@@ 4.2 Graph construction
QIM-based steganography induces significant changes, principally in the correlation of adjacent frame codewords, i.e., edges between codewords of adjacent frames [yang2022real, lin2018rnn, li2012detection, li2017steganalysis, yang2018steganalysis]. Consequently, each of the $T$ frames in the speech signal is represented as a node, forming a graph $G$ composed of $T$ nodes:
@@
  • The edges $e_{j}$, which link the nodes $v_{i}$ and $v_{i+1}$, are directional and symbolize the transitions between two successive frames. This choice of directed edges is crucial for capturing the temporal sequence and dependencies inherent in speech signals, allowing the model to learn how information flows and changes over time, particularly between consecutive frames. This design choice aligns with prior studies in speech processing, where directed relationships are commonly employed to model temporal dependencies and improve learning performance [shirian2021compact, li2023speech].
@@ 4.3.1 GraphSAGE network
During the graph updating phase, the node embeddings within the graph undergo a series of operations aimed at refining and modifying them. This process is crucial for extracting essential features in steganalysis. In our architecture, three GraphSAGE convolution layers are employed to progressively extract more abstract features from the input graph data. The GraphSAGE algorithm, suggested by Hamilton et al. [hamilton2017inductive], is instrumental in this process.
@@
where $h_{u}^{k-1}$ corresponds to the output of node $u$ from the previous layer. The embeddings of all nodes $u$ within the neighbourhood of node $v$ are collectively merged to form the embedding of node $v$ at layer $k$. In Equation 7, various aggregation functions such as mean, pooling, graph convolution, or LSTM can be applied. Within the scope of our model, the LSTM architecture achieved superior performance, as proved later in experiments.
The LSTM aggregator processes the set of neighbor representations as a sequence through an LSTM. For each node $v$:
@@
Figure 5: A two-layer node embedding process. $h_{v}^{0}$ denotes the initial attribute of node $v$. In this illustration, node embedding is shown for the target node $A$. Notably, the embedding of other nodes occurs in parallel.
@@ 4.3.2 Classification task
$\text{Loss}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log\left(\frac{\exp(\hat{y_{i}})}{\sum_{j}\exp(\hat{y_{j}})}\right)\right.\left.+(1-y_{i})\log\left(1-\frac{\exp(\hat{y_{i}})}{\sum_{j}\exp(\hat{y_{j}})}\right)\right]$ (15)
where Loss is the loss value to be minimized, $y_{i}$ is the true label, $\hat{y_{i}}$ are the logits for the $i$-th sample, and $N$ is the number of samples in the dataset. $\sum_{j}\exp(\hat{y_{j}})$ represents the sum of the exponentials of logits across all classes, used for normalization in the softmax function.
@@ 5.1 Experiment setup
The experiments were conducted using a dataset described in [lin2018rnn] and available on the Github [^1] https://github.com/fjxmlzn/RNN-SM platform. This dataset includes 72 hours of English speech and 41 hours of Chinese speech, totalling 320 recordings in 16-bit PCM format. The speech samples were sourced from the internet and included contributions from various male and female speakers. In our methodology, the G.729A LBR speech codec was employed to encode the original speech samples, forming the basis of our cover speech dataset. To create the stego samples, the CNV-QIM [xiao2008approach] steganography method was used, inserting secret bits during the G.729A encoding phase.
@@ 5.2 Evaluation metrics
[…]
@@ 5.3 Detection performance
    A preliminary assessment of the degree of difficulty faced by the proposed GNN scheme in uncovering the CNV-QIM steganographic method, particularly at low embedding rates, was conducted using statistical analysis by comparing cover and stego datasets. Figure 7 presents this analysis using boxplots, contrasting cover and stego datasets across various embedding rates. The results demonstrate that CNV-QIM maintains consistent mean values across different scenarios, with only slight variations in the first (Q1) and third (Q3) quartiles as embedding rates increase. This stability indicates that CNV-QIM is a challenging steganographic method. The evaluation of the proposed steganalysis technique under different embedding rates for 10-second sample segments, as presented in Table 3, reveals nuanced performance characteristics. Notably, the detection capability, as measured across all metrics (Accuracy, Precision, Recall, F1-score), exhibits a slight decreasing trend with lower embedding rates. This is attributed to the fact that higher embedding rates induce more pronounced changes in codewords, providing a more discernible pattern for steganalysis. It is noteworthy that even at a low embedding rate of 20%, our approach maintains high accuracy, registering at 95.17%, and demonstrates strong, balanced performance across the other metrics, with Precision at 95.43%, Recall at 94.95%, F1-score at 95.19%, and AUC at 95.20%.
    To further illustrate this robust performance in the challenging low-embedding-rate scenario, Figure 8 presents the detailed confusion matrix for the 20% embedding rate. Out of a total of 3166 test samples (comprising 1583 actual stego and 1583 actual cover samples), the model successfully identified 1503 TP (correctly classified stego) and 1511 TN (correctly classified cover). Critically, the false classifications were minimal, with only 72 FP (cover incorrectly flagged as stego) and 80 FN (stego samples missed). This granular breakdown underscores the model’s effectiveness in minimizing both false alarms and missed detections, which are crucial for practical steganalysis applications. The proposed steganalysis approach consistently achieves high detection rates despite the inherent challenge posed by the resilience of CNV-QIM, showcasing its strong discriminative power even under subtle embedding.
@@
    Performance results under varying speech segment lengths: In VoIP steganalysis, the length of speech segments is also a crucial factor influencing detection accuracy. A successful approach must achieve a sufficiently high level of accuracy within a limited time detection window. Table 4 explores the impact of sample length on various detection performance metrics, focusing on a fixed embedding rate of 100%. The findings reveal that longer sample lengths lead to a steady and significant improvement across all performance metrics: Accuracy, Precision, Recall, and F1-score. This improvement is rapid, with performance reaching near-perfect levels (e.g., 99.97% accuracy at 3 seconds) and achieving a perfect 100% across all metrics from 7-second samples onwards. This can be attributed to the larger sample length providing more data for steganalysis, allowing the algorithm to learn richer patterns and longer-range dependencies. Crucially, even with short 0.5-second segments, the approach we propose attains exceptionally high detection performance, with Accuracy, Precision, Recall, and F1-score all exceeding 98%. The remarkable consistency across these metrics for each sample length highlights the model’s robust and balanced performance, demonstrating its ability to minimize both false positives and false negatives. The method therefore enables effective detection of QIM steganography even for very short segments of monitored VoIP traffic, which is vital for real-world scenarios where covert data might be dispersed in brief bursts.
@@
    Performance results with varying speech segment lengths and embedding rate: Figure 9 provides the results of detection accuracy with varying embedding rates for sample segments of different lengths. The results reaffirm the method’s reliability, with a general upward trend in accuracy as the embedding rate and sample length increase. It is evident that achieving accurate detection under low embedding rates with short segments poses a challenge, yet our approach maintains an acceptable accuracy exceeding 70%. This challenging scenario represents the Achilles’ heel of our approach, signalling potential areas for further investigation and model enhancements.
@@
To further substantiate the discriminative capabilities of our approach, t-distributed stochastic neighbour embedding (t-SNE) was exploited to visualize the graph-level representation vectors (Equation 12), before and after training, utilizing a set of 2000 graphs selected from the training dataset. Figure 10 presents snapshots of the distribution of graph-level representation vectors at different embedding rates while utilizing a fixed sample length of 0.5 seconds. The colour-coded representation illustrates the gradual separation of stego and cover graphs in the vector space of features post-training. The overlap diminishes as the embedding rate increases, reaching almost complete separation at a 100% embedding rate. These visualizations intuitively showcase the model’s efficacy in extracting and analyzing steganographic speech features across various embedding rates. As a result, the combination of sensitivity to various embedding rates and effectiveness with short samples suggests that our method is versatile and applicable across a wide range of steganographic scenarios, from high-capacity hidden messages to more subtle, security-conscious embeddings. These characteristics position our approach as a powerful tool in the ongoing challenge of audio steganalysis, capable of adapting to different steganographic strategies while maintaining high detection accuracy. These findings provide valuable insights into how the model can be applied in practical VoIP security systems, as discussed in the following section. Moreover, these findings offer valuable insights into the practical applicability of our model in real-world VoIP security systems, which will be further explored in the section 6.
@@ 5.4 Comparison with different model variants
Table 5 presents the results of the experiment, showing the detection accuracy for each architecture variant. The results highlight the significance of various components in the proposed architecture. It is obvious that the complete model (#1\#1# 1) yields the highest level of accuracy, demonstrating the effectiveness of our GraphSAGE for steganalysis. Variant #2\#2# 2, which removes the first and second max-pooling layers, experiences a significant drop in accuracy, emphasizing the importance of using the hierarchical pooling approach in capturing fine-grained details and high-level patterns. This is particularly beneficial in scenarios with low embedding rates where changes produced by steganography are minimal. In variant #3\#3# 3, the use of a mean aggregator instead of LSTM highlights the importance of the latter in our approach. The LSTM aggregator captures sequential dependencies and temporal patterns in the input data, which is crucial for steganalysis tasks dealing with speech segments. By leveraging LSTM, the model can effectively learn inter-frame relationships and sequential context, allowing for more effective detection of subtle changes introduced by QIM steganography. Variant #4\#4# 4, using max pooling instead of mean pooling, also leads to an accuracy of 85%, significantly inferior to the complete model. A possible explanation is that max-pooling selects the maximum value from each layer, discarding the information about the overall distribution of features within the layer. In the context of QIM steganalysis, where capturing collective characteristics is crucial for effective detection, this loss of information makes it harder for the model to notice subtle patterns that indicate steganographic content. Mean pooling, on the other hand, averages the features at each layer, providing a more comprehensive representation of the layer’s characteristics and proving to be more effective for our specific steganalysis task. Furthermore, variant #5\#5# 5, where GraphSAGE layers are replaced with GCN, exhibits a notable decrease in accuracy, suggesting that the specific architecture of GraphSAGE contributes significantly to the model’s performance. Similarly, when GraphSAGE layers are replaced with graph attention network (GAT) layers (Variant #6\#6# 6), while performance is strong (91.63%), it remains lower than our proposed GraphSAGE model. This indicates that for this specific VoIP steganalysis task, the fixed aggregation approach of GraphSAGE, which comprehensively samples and aggregates features from a fixed neighborhood, provides a more effective representation for QIM-induced subtle changes than the attention-based weighting of GAT layers. Finally, in variants #7\#7# 7 and #8\#8# 8, the first reduces the number of GraphSAGE layers by one, while the second adds one extra GraphSAGE layer. Both configurations result in a decrease in terms of accuracy, consequently supporting the suitability of employing three layers in our model architecture. Generally, a deeper network has the potential to capture more intricate features within the input data. Variant #8\#8# 8’s performance highlights a crucial observation – the depth of the network does not necessarily correlate with improved model performance. Deeper networks, as seen in variant #8\#8# 8, may face challenges such as overfitting and vanishing gradient problems, emphasizing the importance of striking a balance between network depth and effective steganalysis. The optimal configuration should not only capture relevant features but also mitigate potential issues associated with increased model complexity.

### v2

@@ 3.2 QIM-based Steganography
$\left\{\begin{aligned} &C_{i}=C^{1}_{i}\cup C^{2}_{i}\\ &C^{1}_{i}\cap C^{2}_{i}=\emptyset\end{aligned}\right$. (2)
@@
The challenge in QIM steganography arises from increased distortion resulting from the use of a reduced codeword set in the quantization procedure. The fundamental issue lies in how to partition the original codebook $C_{i}$ into multiple sub-codebooks. Researchers have proposed various methods including matrix encoding QIM (ME-QIM) [10], complementary neighbour vertices QIM (CNV-QIM) [57], and others.
@@ 3.3 GNN Overview
Here, $h_{v}^{k}$ and $h_{v}^{k-1}$ represent the hidden states of node $v$ at layers $k$ and k1k-1, respectively. The neighborhood of a node $v$ is symbolized as $N(v)$, and $w_{uv}^{k-1}$ represents the weight associated to the edge connecting $u$ and $v$ in the (k1)(k{-}1)th layer. The function $\mathcal{F}$ is a non-linear activation function. The update rule essentially aggregates information from neighbouring nodes to refine the current node’s hidden state.
A widely adopted family of GNN models includes graph convolutional networks (GCNs) [58], gated graph neural networks (GGNNs) [59], graph sample and aggregation (GraphSAGE) [60], and more. These architectures often involve variations of the basic propagation rule to capture more complex relationships within the graph.
@@ 4 Proposed framework
Building upon the foundational concepts of speech coding, QIM steganography, and GNN discussed in Section 3, we now present the detailed architecture and components of our proposed steganalysis system. Specifically, this section describes an efficient method designed to uncover QIM steganography within compressed, LBR VoIP streams. Our approach is specifically tailored to work with speech compressed using the G.729A codec, which is widely used in VoIP communications. Figure 2 illustrates the structure of our proposed VoIP steganalysis, while Algorithm 1 outlines the steps involved.
As depicted in the figure, our framework consists of a three-stage pipeline. First, the Graph Construction Module takes a raw VoIP stream, extracts its Quantization Index Sequence (QIS) matrix, and converts it into a directed acyclic graph representation. Second, this graph is processed by the GraphSAGE Network Module, which learns a discriminative graph-level feature vector through hierarchical message passing and pooling operations. Finally, the resulting feature vector is fed into the Classification Network Module, which performs a binary classification to determine if the stream is ’Cover’ (benign) or ’Stego’ (contains hidden data). The remainder of this section provides a detailed explanation of each of these components, covering codewords correlation analysis, graph construction, and the network architecture.
Figure 2: The overall structure of the proposed approach, which consists of a three-stage pipeline. (1) Graph Construction Module: A QIS matrix is extracted from the input VoIP stream and converted into a directed acyclic graph. (2) GraphSAGE Network Module: The graph is processed via graph updating and pooling to generate a representative feature vector. (3) Classification Network Module: The final vector is passed through dropout and linear layers to classify the stream as ’Cover’ or ’Stego’.
@@
13:Construct adjacency matrix AA:
@@ 4.1 Codewords correlation analysis
Due to the repeating patterns in human speech sounds (especially vowels and voiced consonants), the vocal signal exhibits stability over limited durations, approximately the frame length. Consequently, the codewords exhibit correlations within the same frame. Furthermore, speech signals have local periodicity, meaning that codewords have similar values in different frames. As a result, there are four types of inter-codeword correlations, namely: intra-frame correlation, successive frame correlation, cross-frame correlation, and cross-word correlation [16]. Local features are characterized by the first two correlations, while global ones are reflected by the latter two.
The QIM steganography process has a direct impact on the correlations between these codewords, thereby altering their statistical distribution. Earlier VoIP steganalysis approaches aimed to extract correlation features to uncover steganograms, as in [37, 39].
@@ 4.2 Graph construction
QIM-based steganography induces significant changes, principally in the correlation of adjacent frame codewords, i.e., edges between codewords of adjacent frames [14, 16, 36, 37, 39]. Consequently, each of the $T$ frames in the speech signal is represented as a node, forming a graph $G$ composed of $T$ nodes:
@@
  • The edges $e_{j}$, which link the nodes $v_{i}$ and $v_{i+1}$, are directional and symbolize the transitions between two successive frames. This choice of directed edges is crucial for capturing the temporal sequence and dependencies inherent in speech signals, allowing the model to learn how information flows and changes over time, particularly between consecutive frames. This design choice aligns with prior studies in speech processing, where directed relationships are commonly employed to model temporal dependencies and improve learning performance [61, 62].
@@ 4.3.1 GraphSAGE network
During the graph updating phase, the node embeddings within the graph undergo a series of operations aimed at refining and modifying them. This process is crucial for extracting essential features in steganalysis. In our architecture, three GraphSAGE convolution layers are employed to progressively extract more abstract features from the input graph data. The GraphSAGE algorithm, suggested by Hamilton et al. [60], is instrumental in this process.
@@
where $h_{u}^{k-1}$ corresponds to the output of node $u$ from the previous layer. The embeddings of all nodes $u$ within the neighbourhood of node $v$ are collectively merged to form the embedding of node $v$ at layer $k$. In Equation 7, various aggregation functions such as mean, pooling, graph convolution, or LSTM can be applied. Within the scope of our model, the LSTM architecture achieved superior performance, as proved later in experiments. The LSTM aggregator processes the set of neighbor representations as a sequence through an LSTM. For each node $v$:
@@
Figure 5: A two-layer node embedding process. $h_{v}^{0}$ denotes the initial attribute of node $v$. In this illustration, node embedding is shown for the target node AA. Notably, the embedding of other nodes occurs in parallel.
@@ 4.3.2 Classification task
[…]
@@ 5.1 Experiment setup
The experiments were conducted using a dataset described in [16] and available on the Github [^1] https://github.com/fjxmlzn/RNN-SM platform. This dataset includes 72 hours of English speech and 41 hours of Chinese speech, totalling 320 recordings in 16-bit PCM format. The speech samples were sourced from the internet and included contributions from various male and female speakers. In our methodology, the G.729A LBR speech codec was employed to encode the original speech samples, forming the basis of our cover speech dataset. To create the stego samples, the CNV-QIM [57] steganography method was used, inserting secret bits during the G.729A encoding phase.
@@ 5.2 Evaluation metrics
$\text{F1-score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision + Recall}}=\frac{2\times\text{TP}}{2\times\text{TP + FP + FN}}$ (19)
@@ 5.3 Detection performance
    A preliminary assessment of the degree of difficulty faced by the proposed GNN scheme in uncovering the CNV-QIM steganographic method, particularly at low embedding rates, was conducted using statistical analysis by comparing cover and stego datasets. Figure 7 presents this analysis using boxplots, contrasting cover and stego datasets across various embedding rates. The results demonstrate that CNV-QIM maintains consistent mean values across different scenarios, with only slight variations in the first (Q1) and third (Q3) quartiles as embedding rates increase. This stability indicates that CNV-QIM is a challenging steganographic method. The evaluation of the proposed steganalysis technique under different embedding rates for 10-second sample segments, as presented in Table 4, reveals nuanced performance characteristics. Notably, the detection capability, as measured across all metrics (Accuracy, Precision, Recall, F1-score), exhibits a slight decreasing trend with lower embedding rates. This is attributed to the fact that higher embedding rates induce more pronounced changes in codewords, providing a more discernible pattern for steganalysis. It is noteworthy that even at a low embedding rate of 20%, our approach maintains high accuracy, registering at 95.17%, and demonstrates strong, balanced performance across the other metrics, with Precision at 95.43%, Recall at 94.95%, F1-score at 95.19%, and AUC at 95.20%. This high level of accuracy in a low-payload scenario is particularly significant, as attackers frequently use low embedding rates to minimize statistical disturbances and evade detection. The model’s ability to maintain such robust performance demonstrates its high sensitivity to even the most subtle steganographic artifacts.
    To further illustrate this robust performance in the challenging low-embedding-rate scenario, Figure 8 presents the detailed confusion matrix for the 20% embedding rate. Out of a total of 3166 test samples (comprising 1583 actual stego and 1583 actual cover samples), the model successfully identified 1503 TP (correctly classified stego) and 1511 TN (correctly classified cover). Critically, the false classifications were minimal, with only 72 FP (cover incorrectly flagged as stego) and 80 FN (stego samples missed). This granular breakdown is crucial for practical applications, as it demonstrates the model’s ability to strike a critical balance: minimizing false alarms (FP), which could disrupt legitimate communication, while simultaneously minimizing missed detections (FN), which would allow covert channels to go unnoticed. The proposed steganalysis approach consistently achieves high detection rates despite the inherent challenge posed by the resilience of CNV-QIM, showcasing its strong discriminative power even under subtle embedding. This demonstrates that the GNN-based architecture successfully learns to distinguish the subtle, distributed changes introduced by steganography from the natural variations in cover speech.
@@
    Performance results under varying speech segment lengths: In VoIP steganalysis, the length of speech segments is also a crucial factor influencing detection accuracy. A successful approach must achieve a sufficiently high level of accuracy within a limited time detection window. Table 5 explores the impact of sample length on various detection performance metrics, focusing on a fixed embedding rate of 100%. The findings reveal that longer sample lengths lead to a steady and significant improvement across all performance metrics: Accuracy, Precision, Recall, and F1-score. This improvement is rapid, with performance reaching near-perfect levels (e.g., 99.97% accuracy at 3 seconds) and achieving a perfect 100% across all metrics from 7-second samples onwards. This trend is directly linked to the nature of our GNN-based model; a longer speech segment translates to a larger graph with more nodes and edges. This allows the GraphSAGE architecture to aggregate information over a wider neighborhood, resulting in more stable and representative graph-level embeddings that make the distinction between cover and stego patterns far more pronounced. Crucially, even with short 0.5-second segments, the approach we propose attains exceptionally high detection performance, with Accuracy, Precision, Recall, and F1-score all exceeding 98%. This is a critical finding for practical deployment, as real-time monitoring systems must often make decisions based on individual or fragmented data packets. Furthermore, attackers may intentionally disperse hidden data in short, disjointed bursts to evade detection systems that require longer samples for analysis. The remarkable consistency across these metrics for each sample length highlights the model’s robust and balanced performance, demonstrating its ability to minimize both false positives and false negatives. The method therefore enables effective detection of QIM steganography even for very short segments of monitored VoIP traffic, which is vital for real-world scenarios where covert data might be dispersed in brief bursts.
@@
    Performance results with varying speech segment lengths and embedding rate: Figure 9 provides the results of detection accuracy with varying embedding rates for sample segments of different lengths. The results reaffirm the method’s reliability, with a general upward trend in accuracy as the embedding rate and sample length increase. This trend is expected, as longer samples and higher embedding rates provide the model with more data and a stronger, less ambiguous signal, respectively, making the detection task easier. It is evident that achieving accurate detection under low embedding rates with short segments poses a challenge for any steganalysis system. This scenario, representing the most difficult detection conditions due to minimal available information, can be considered a benchmark for model robustness. Our approach maintains an acceptable accuracy exceeding 70% even in these worst-case conditions, which, rather than being an ”Achilles’ heel,” demonstrates that the GNN is still able to capture a discriminative signal from extremely sparse and noisy data, highlighting its effectiveness.
@@
To further substantiate the discriminative capabilities of our approach, t-distributed stochastic neighbour embedding (t-SNE) was exploited to visualize the graph-level representation vectors (Equation 12), before and after training, utilizing a set of 2000 graphs selected from the training dataset. Figure 10 presents snapshots of the distribution of graph-level representation vectors at different embedding rates while utilizing a fixed sample length of 0.5 seconds. Before training (top-left), the cover and stego data points are heavily overlapped, indicating that they are indistinguishable in their raw feature space. The colour-coded representation illustrates the gradual separation of stego and cover graphs in the vector space of features post-training. The overlap diminishes as the embedding rate increases, reaching almost complete separation at a 100% embedding rate. This directly correlates the strength of the steganographic signal with the distance between the clusters in the learned space, visually confirming the model’s sensitivity. These visualizations intuitively showcase the model’s efficacy in extracting and analyzing steganographic speech features across various embedding rates. As a result, the combination of sensitivity to various embedding rates and effectiveness with short samples suggests that our method is versatile and applicable across a wide range of steganographic scenarios, from high-capacity hidden messages to more subtle, security-conscious embeddings. These characteristics position our approach as a powerful tool in the ongoing challenge of audio steganalysis, capable of adapting to different steganographic strategies while maintaining high detection accuracy. These findings provide valuable insights into how the model can be applied in practical VoIP security systems, as discussed in the following section. Moreover, these findings offer valuable insights into the practical applicability of our model in real-world VoIP security systems, which will be further explored in the section 6.
@@ 5.4 Comparison with different model variants
Table 6 presents the results of the experiment, showing the detection accuracy for each architecture variant. The results highlight the significance of various components in the proposed architecture. It is obvious that the complete model (#1\#1) yields the highest level of accuracy, demonstrating the effectiveness of our GraphSAGE for steganalysis. Variant #2\#2, which removes the first and second max-pooling layers, experiences a significant drop in accuracy, emphasizing the importance of using the hierarchical pooling approach in capturing fine-grained details and high-level patterns. This is particularly beneficial in scenarios with low embedding rates where changes produced by steganography are minimal. In variant #3\#3, the use of a mean aggregator instead of LSTM highlights the importance of the latter in our approach. The LSTM aggregator captures sequential dependencies and temporal patterns in the input data, which is crucial for steganalysis tasks dealing with speech segments. By leveraging LSTM, the model can effectively learn inter-frame relationships and sequential context, allowing for more effective detection of subtle changes introduced by QIM steganography. Variant #4\#4, using max pooling instead of mean pooling, also leads to an accuracy of 85%, significantly inferior to the complete model. A possible explanation is that max-pooling selects the maximum value from each layer, discarding the information about the overall distribution of features within the layer. In the context of QIM steganalysis, where capturing collective characteristics is crucial for effective detection, this loss of information makes it harder for the model to notice subtle patterns that indicate steganographic content. Mean pooling, on the other hand, averages the features at each layer, providing a more comprehensive representation of the layer’s characteristics and proving to be more effective for our specific steganalysis task. Furthermore, variant #5\#5, where GraphSAGE layers are replaced with GCN, exhibits a notable decrease in accuracy, suggesting that the specific architecture of GraphSAGE contributes significantly to the model’s performance. Similarly, when GraphSAGE layers are replaced with graph attention network (GAT) layers (Variant #6\#6), while performance is strong (91.63%), it remains lower than our proposed GraphSAGE model. This indicates that for this specific VoIP steganalysis task, the fixed aggregation approach of GraphSAGE, which comprehensively samples and aggregates features from a fixed neighborhood, provides a more effective representation for QIM-induced subtle changes than the attention-based weighting of GAT layers. Finally, in variants #7\#7 and #8\#8, the first reduces the number of GraphSAGE layers by one, while the second adds one extra GraphSAGE layer. Both configurations result in a decrease in terms of accuracy, consequently supporting the suitability of employing three layers in our model architecture. Generally, a deeper network has the potential to capture more intricate features within the input data. Variant #8\#8’s performance highlights a crucial observation – the depth of the network does not necessarily correlate with improved model performance. Deeper networks, as seen in variant #8\#8, may face challenges such as overfitting and vanishing gradient problems, emphasizing the importance of striking a balance between network depth and effective steganalysis. The optimal configuration should not only capture relevant features but also mitigate potential issues associated with increased model complexity.